1. A machine gun fires bullets, each with mass m = 55 g at a speed of u = 450 m.s^−1. The gun fires 5.0 bullets per second. Let's use it as a rocket engine. What thrust does it produce?

Thrust _____ N.

2. A mass breaks suddenly into two parts, masses M and m, with speeds V and v respectively and carrying a total kinetic energy K. If M / m = R = 4.2, what fractions of K does each of the masses carry? (Think about significant figures.)

a) M carries _____ K.

b) m carries _____ K.

Please separate your answers with a comma. For example: if you think that each carries half of the total kinetic energy, write 0.50, 0.50.
1
Expert's answer
2020-03-13T10:50:13-0400
  1. As per the given question,

mass of the bullets=55gm =0.055kg

speed of the bullets=450 m/sec

Number of bullets fired per second(n) =5

we know that force = rate of change of momentum

So, thrust force "F=\dfrac{nmv}{t}"

"\Rightarrow F=\dfrac{5\times 0.055}{1} =0.275N"